OpenStudy (anonymous):

A principal of 835 dollars i invested in an account at 2 percent per quarter simpleinterest. which of the following squencs describe th dollar amount of principal in the account at quarterly intervals. a.16.70, 33.40, 50.10, 66.80, 83.50,... b.835.00, 901.80, 968.60, 1035.40, 1102.20,... c.835.00, 851.70, 868.40, 885.10, 901.80,.... d.66.80, 133.60, 200.40, 267.20, 334.00 e.83.00, 851.70, 868.73, 886.11, 903.83,...

OpenStudy (campbell_st):

start with 835 1 quarter I = 835*0.02*1 = 16.70 + 835 = 851.70 2nd quarter I =851.70*0.02*1 = 17.03 + 851.70 = 868.73 3rd quarter I = 868.73*0.02*1 = 17.37 + 868.73 = 886.10 so perhaps the last choice

OpenStudy (radar):

The first entry in the last choice is $83.00...........where did that come from? I don't agree with the last choice.

OpenStudy (radar):

Unless it is a type and was supposed to be $835.00 which is the initial deposit?????